OpenStudy (anonymous):

Find the distance between the following pairs of points: (-4,3),(2,5)

OpenStudy (anonymous):

\[d=\sqrt{(-4-2)^2+(5-3)^2}\]

OpenStudy (akashdeepdeb):

You know everything about this question! What seems to be the problem here @Algebra2013 ? :D

OpenStudy (anonymous):

I got \[\sqrt{40} for an answer but \it is wrong???\]

OpenStudy (anonymous):

I came up with 40 squared which is wrong? please help

OpenStudy (akashdeepdeb):

Are you sure those ARE the points? Because sqrt 40 IS THE RIGHT ANSWER! :D

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

correct points answers are::::: \[2\sqrt{10} , \sqrt{28} , \sqrt{8} , or 10\]

OpenStudy (anonymous):

Anyone come up with one of these??

OpenStudy (akashdeepdeb):

2 sqrt 10 is correct! :) Because sqrt 40 can be written as \[\sqrt{2*2*10} = 2 * \sqrt{10}\] Understood? :)

OpenStudy (anonymous):

Yes, thank you for the much needed help

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

Mary and Jane are both at the gas station at Hill and Elm. From the station, Mary drives 10 miles east and Jane drives 15 miles south. Find the distance between Mary and Jane

OpenStudy (akashdeepdeb):

Do you know Pythagorus' theorem?

OpenStudy (anonymous):

midpoint??

OpenStudy (anonymous):

\[=(\frac{ x+X }{ 2 }),(\frac{ y+y }{ 2 })=\]

OpenStudy (akashdeepdeb):

No No Pythagorus' theorem? Ever heard of it or not? :/

OpenStudy (anonymous):

No

OpenStudy (anonymous):

is there a formula?

OpenStudy (anonymous):

Did you get 5 for an answer?

OpenStudy (akashdeepdeb):

I suggest you go through this before the formula! :) http://www.mathwarehouse.com/geometry/triangles/how-to-use-the-pythagorean-theorem.php

OpenStudy (anonymous):

ok there are only 2 numbers we can use which are 10 miles east and 15 miles south

OpenStudy (akashdeepdeb):

Yes! And they form a right angle inbetween them. :) Read the page I sent you.

OpenStudy (anonymous):

\[10^{2}+15^{2}=X\]

OpenStudy (akashdeepdeb):

10^2+15^2=X^2

OpenStudy (akashdeepdeb):

Not X X^2 :D

OpenStudy (anonymous):

I got 18

OpenStudy (anonymous):

=325 (18)(18)=325 x=18

OpenStudy (anonymous):

\[5\sqrt{13}, -5\sqrt{13}\]

OpenStudy (anonymous):

got it thank you very much

OpenStudy (akashdeepdeb):

Yes nearly very close to 18!! :D :)

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

if the midpoint of two points is (5,-1) and one of the points is (4,-3), find the other point

OpenStudy (akashdeepdeb):

Just use the formula you told me! And take the other point as say (x,y) Understood? :)

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

9 doesn't go into 2????

OpenStudy (anonymous):

\[\frac{ 5+4 }{ 2}, \frac{ -1+-3 }{ 2 }\]

OpenStudy (akashdeepdeb):

See take it like you told me yourself \[Midpoint = (\frac{ x1 + x2}{ 2 } , \frac{ y1 + y2 }{ 2 })\] Now let the other point be = (x,y) So we can also write the above equation as \[(5,-1) = (\frac{ 4 + x}{ 2 } , \frac{ -3 + y }{ 2 })\] So \[5 = \frac{ 4+x }{ 2 } AND -1 = \frac{ -3+y }{ 2 }\]\

OpenStudy (anonymous):

x=6, y=1

OpenStudy (anonymous):

(6,1)

OpenStudy (akashdeepdeb):

Yes! You got it! :D

OpenStudy (anonymous):

Thank you so much!!!!!!!!!!!

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

A painter is going to paint the trim of a window located 18 feet from the ground. If the base of the ladder is placed 9 feet from the side of the house, how long does the ladder need to be to reach the window?

OpenStudy (akashdeepdeb):

Use Pythagorus here too. See @Algebra2013 I cannot help you with all questions. You have to figure most out yourself. :)